OpenStudy (anonymous):

One year after a celebrity firs began posting comments on a website, she had 20,500 followers. Three years after this celebrity first began posting comments, she had 93,500 followers. A) Find the slope of the line that passes through the points (1,20500) and (3,93500) B) interpret the slope as a rate of change

OpenStudy (anonymous):

Hahahahhahaha!!! I'm from India and even though I know how to solve it I think Callisto will guide you through!! But boiy I wish i had questions like these when I was in school!! I would had started to love maths then itself.. :D

OpenStudy (callisto):

A) slope = \(\frac{y_2-y_1}{x_2-x_1}\) E.G: take (1,20500) as (\(x_1, y_1\)) What is (\(x_2, y_2\))? B) Rate of change = slope = ... followers/year

OpenStudy (anonymous):

93500-20500 x2 and x1 = 3-1

OpenStudy (anonymous):

rate is 36,500?

OpenStudy (anonymous):

The rate over three year period is 12,166

OpenStudy (anonymous):

slope is 36,500